OpenStudy (anonymous):

Solve the following quadratic equation by a method of your choice: x^2 + 3x = 0. A. x = 0 or -3 B. x = 0 or 4 C. x = 1 or 3 D. x = -3 or 4

OpenStudy (anonymous):

A

OpenStudy (anonymous):

A

sam (.sam.):

x^(2)+3x=0 x(x+3)=0 x=0 x+3=0 x=-3 x=0,-3